Suppose that when your income increases by $300, your consumption expenditures increases by $225.

Your marginal propensity to consume (MPC) is _________ .

If your MPC was the same as the MPC for the economy as a whole, the expenditure multiplier for the economy would be ______________ .

Thus, a $2 million investment project would increase income by $ _________ million in total.
Answer #1

Answer

MPC=change in consumption/change in income

=225/300

MPC=0.75

the multiplier =1/(1-MPC)

=1/(1-0.75)=4

the multiplier is 4

==========

change in income =change in investment * multiplier

=2*4

=$8 million

the income increases by maximum of $8 million.